# Constrained dynamics and confinement in the two-dimensional quantum Ising model

### v2

@@ Acknowledgements.
We acknowledge enlightening discussions with Marcello Dalmonte, Roberto Verdel, Simone Notarnicola, Jernej Rudi Finžgar, and Sourav Nandy. The authors acknowledge funding from CN00000013 - Italian Research Center on HPC, Big Data and Quantum Computing (ICSC), the H2020 projects EuRyQa, QuantERA2017 project QuantHEP, QuantERA2021 project T-NiSQ, the Quantum Technology Flagship project PASQuanS2, the Italian Ministry of University and Research (MUR) via PRIN2022 project TANQU, the Departments of Excellence grant 2023-2027 Quantum Frontiers, the German Federal Ministry of Education and Research (BMBF) the funding program quantum technologies, project QRydDemo, and the World Class Research Infrastructure - Quantum Computing and Simulation Center (QCSC) of Padova University. The simulations were performed using the Quantum Green Tea software version 0.3.23 and Quantum Tea Leaves version 0.4.46 [48]. The authors acknowledge computational resources of the HPC center Vega at the Institute of Information Science (IZUM) in Maribor, Slovenia and HPC Leonardo by Cineca, Italy.
Supplementary Material
@@
The numerical calculations presented in this work are performed with tree tensor networks (TTNs) [28, 29]. In this appendix, we provide qualitative reasoning for the utility of TTNs compared to other tensor networks and technical aspects of the algorithm for time evolution that we apply, and present plots of magnetization and entanglement obtained for a set of maximal bond dimensions.
@@
In TTNs, the state is represented by a tensor network in the shape of a symmetric binary tree, where the bottom tensors (’leaves’ of the tree) carry the physical indices. The tree tensor network ansatz is particularly successful at describing states where distant sites are entangled, such as in chains with long-range interactions. The reason why can be understood through a qualitative picture of the way quantum states are encoded in tensor networks. Qualitatively, entanglement between two correlated physical states is transferred through bonds of the network that connect them. The more entanglement there is, the bigger the bond dimension required to accurately describe the state. Generally, this also means entangled states are better represented by tensor networks with a bigger number of links. An extreme example of this is the projected entangled-pair state (PEPS) ansatz, where the tensor network mimics the structure of the 2D lattice and each tensor is directly connected to its nearest neighbours. However, this introduces loops in the network, which leads to technical difficulties. Namely, it is not possible to define the isometry center of a network with loops. Therefore, the algorithms for contracting such networks cannot rely on the underlying symmetries. This leads to poor scaling of calculations of local observables with bond dimension, and limits PEPS to bond dimensions on the order of $\chi\sim 10$. The TTN ansatz is in between the MPS and PEPS in the sense that it is the best connected tensor network which by construction has no loops. Furthermore, because two physical sites separated by $m$ lattice sites in real space are only separated by $\sim\ln m$ in the TTN, the entanglement between them is captured more efficiently.
@@
The most common way to evolve tensor network states in time is with the time-dependent variatonal principle (TDVP) [31, 32]. The most popular is the two-tensor version of the algorithm (TDVP2), where at each step two neighbouring tensors are contracted, evolved in time, and then split and truncated via SVD. This algorithm can grow the bond dimensions between tensors at each time step and is thus a natural choice for simulations starting from weakly entangled states, such as the product states we use. A much faster version of the same concept is the single-tensor TDVP (TDVP1), where each tensor is optimized without contractions with the neighbours. In this case, the bond dimension cannot change, which presents a problem if the initial state is a product state with $\chi=1$.
However, we find that the calculations are much faster if one pads the initial product state with numerical noise on the order of $10^{-16}$ up to bond dimension $\chi$, and we perform the evolution with TDVP1, without truncation. We find a speedup in computational time of about 30×30$\times$30 × for a $8\times 8$ system at $\chi=100$. We also tried hybrid approaches (like performing TDVP2 every $n-$th time step and TDVP1 otherwise, with $n$ around 10). The advantage here is that the initial bond dimension is not $\chi$ but rather gradually increases every $n$-th step. This speeds up the initial time steps, but the advantage is quickly lost once the bond dimension reaches a sizeable fraction of the one used for TDVP1. An advantage might be expected for the easiest cases (small $g/J$ limit) where the bond dimension never reaches the maximal allowed value. However, these cases can also be accurately simulated by a TDVP1 simulation with smaller $\chi$, which is anyway much faster. Furthermore, on a technical level having a single convergence parameter ($\chi$) is more desirable compared to having two ($\chi$ and the minimal truncation $\epsilon$) with non-trivial interplay.
@@
In Fig. 5, we show the convergence of (a) magnetization and (b) entanglement with increasing maximal bond dimension $\chi$. The deviation of the magnetization at given $\chi$ with respect with the results obtained with $\chi=160$ are shown in Fig. 5(c).
The convergence is strongly dependent on the underlying physics. Deep in the confined phase at $g\ll J$, results are well converged already for $\chi=60$ or less. At larger $g$, larger $\chi$ is required to obtain well-converged results and the simulations become unreliable much sooner in simulated time.
@@
Figure 5: Convergence with bond dimension $\chi$. (a) Magnetization $\sigma^{z}$ of a single spin calculated with increasing bond dimension $\chi$ for a set of $g/J$. (b) Entanglement entropy calculated with increasing bond dimension $\chi$ for a set of $g/J$. (c) Log scale plots of the deviation of magnetization from the results obtained with $\chi=160$, $\delta\left\langle\sigma^{z}\right\rangle=\left|\left\langle\sigma^{z}\right\rangle(\chi)-\left\langle\sigma^{z}\right\rangle(\chi=160)\right|$.
@@
$H_{0}=-J\sum_{\langle i,j\rangle}\sigma^{z}_{i}\sigma^{z}_{j},$ (3)
@@
$\delta E^{(1)}=-\left\langle\psi\right|V\left|\psi\right\rangle.$ (5)
@@
$\begin{split}\delta E^{(2)}&=-\sum_{k\neq\psi}\left\langle\psi\right|V\left|k\right\rangle\frac{1}{E_{k}-E_{0}}\left\langle k\right|V\left|\psi\right\rangle\\ &=-\sum_{k\neq\psi}\frac{1}{E_{k}-E_{0}}\left|\left\langle\psi\right|V\left|k\right\rangle\right|^{2},\end{split}$ (6)
@@
$\begin{split}\delta E_{0}^{(2)}&=-\sum_{k}\left\langle\psi_{0}\right|V\left|k\right\rangle\frac{1}{E_{k}-E_{0}}\left\langle k\right|V\left|\psi_{0}\right\rangle\\ &=-\sum_{k}\frac{1}{E_{k}-E_{0}}\left|\left\langle\psi_{0}\right|V\left|k\right\rangle\right|^{2}.\end{split}$ (7)
@@
The sum in Eq. (7) thus gives
$\delta E_{0}^{(2)}=-\sum_{i}^{N}\frac{g^{2}}{8J}=-\frac{g^{2}}{8J}N^{2}.$ (8)
@@
$\left|\psi_{1}\right\rangle=\frac{1}{\sqrt{N^{2}}}\sum_{i}\sigma^{x}_{i}\left|\psi_{0}\right\rangle.$ (9)
@@
$\delta E_{1}^{(2)}=\sum_{k}\frac{1}{E_{k}-E_{1}}\left|\left\langle\psi_{1}\right|V\left|k\right\rangle\right|^{2},$ (10)
@@
$\left|k\right\rangle=\sigma^{x}_{i}\sigma^{x}_{j}\left|\psi_{0}\right\rangle.$ (11)
@@
$\begin{split}\left\langle\psi_{0}\right|V\left|\psi_{1}\right\rangle&=\left\langle\psi_{0}\right|-g\sum_{i}\sigma^{x}_{i}\frac{1}{\sqrt{N^{2}}}\sum_{j}\sigma^{x}_{j}\left|\psi_{0}\right\rangle\\ &=-\frac{g}{\sqrt{N^{2}}}\sum_{ij}\left\langle\psi_{0}\right|\sigma^{x}_{i}\sigma^{x}_{j}\left|\psi_{0}\right\rangle\\ &=-\frac{g}{\sqrt{N^{2}}}\sum_{ij}\delta_{ij}\\ &=-\frac{g}{\sqrt{N^{2}}}N^{2},\end{split}$ (12)
@@
$\frac{1}{E_{0}-E_{1}}\left|\left\langle\psi_{0}\right|V\left|\psi_{1}\right\rangle\right|^{2}=-\frac{1}{8J}\times g^{2}N^{2}.$ (13)
@@
$\begin{split}\left\langle\psi_{2}^{\prime}\right|_{kl}V\left|\psi_{1}\right\rangle&=-\frac{g}{\sqrt{N^{2}}}\left\langle\psi_{2}^{\prime}\right|_{kl}\sum_{ij}\sigma^{x}_{i}\sigma^{x}_{j}\left|\psi_{0}\right\rangle\\ &=-\frac{g}{\sqrt{N^{2}}}\times 2,\end{split}$ (14)
@@
$2N^{2}\frac{1}{E_{2}^{\prime}-E_{1}}\left|\left\langle\psi_{2}^{\prime}\right|_{kl}V\left|\psi_{1}\right\rangle\right|^{2}=2N^{2}\times\frac{4g^{2}}{N^{2}}\times\frac{1}{4J}=\frac{2g^{2}}{J}.$ (15)
@@
$\left\langle\psi_{2}^{\prime\prime}\right|_{kl}V\left|\psi_{1}\right\rangle=-\frac{g}{\sqrt{N^{2}}}\times 2,$ (16)
@@
$\frac{1}{2}N^{2}(N^{2}-5)\times\frac{4g^{2}}{N^{2}}\times\frac{1}{8J}=2\frac{g^{2}}{8J}\left(N^{2}-5\right).$ (17)
@@
$\begin{split}\delta E_{1}^{(2)}&=-\left(-\frac{g^{2}}{8J}N^{2}+\frac{2g^{2}}{J}+\frac{g^{2}}{8J}2\left(N^{2}-5\right)\right)\\ &=-\frac{g^{2}}{8J}\left(N^{2}+6\right).\end{split}$ (18)
Importantly, the extensive ($\propto N^{2}$) contribution is the same as for the ground state, see Eq. (8), resulting in a $N$-independent gap.
@@
$\left|\psi_{2}^{\prime}\right\rangle=\frac{1}{2N^{2}}\sum_{ij}\sigma^{x}_{i}\sigma^{x}_{j}\left|\psi_{0}\right\rangle,$ (19)
@@
$\left\langle\psi_{1}\right|V\left|\psi_{2}^{\prime}\right\rangle=\left\langle\psi_{0}\right|\sigma^{x}_{i}\sum_{j}g\sigma^{x}_{j}\frac{1}{\sqrt{2N^{2}}}\sum_{kl}\sigma^{x}_{k}\sigma^{x}_{l}\left|\psi_{0}\right\rangle=g\frac{4}{\sqrt{2N^{2}}}.$ (20)
@@
$N^{2}\times\frac{-1}{4J}\times\frac{16g^{2}}{2N^{2}}.$ (21)
@@
$\begin{split}\left\langle\psi_{3}^{\prime}\right|V\left|\psi_{2}^{\prime}\right\rangle=g\frac{2}{\sqrt{2N^{2}}}.\end{split}$ (22)
@@
$6N^{2}\times\frac{1}{4J}\times\frac{4g^{2}}{2N^{2}}.$ (23)
@@
$\left\langle\psi_{3}^{\prime\prime}\right|V\left|\psi_{2}^{\prime}\right\rangle=g\frac{1}{2N^{2}},$ (24)
@@
$N^{2}\left(N^{2}-8\right)\times\frac{1}{4J}\times\frac{g^{2}}{2N^{2}}.$ (25)
@@
$\delta E_{2}^{(2)}=-g\frac{g^{2}}{8J}N^{2}.$ (26)
See the main text around Fig. 2 for a qualitative interpretation of the obtained result.
@@
In this section, we provide additional results for the dynamics of the square. Fig. 6 shows a grid of magnetization for the lower right quadrant of the square (the panels with blue frames) and two surrounding rows and columns (red frames). Time is rescaled by $g$, as in Fig. 4.
As discussed in the main text, we observe a crossover between two limiting behaviours; at small $g/J$ the dynamics consists of a sequence of resonant spin flips which originate from the corner, while with increasing $g/J$ the spins at the interface melt simultaneously. Spins within the square remain inert until the wave of melting reaches them. Interestingly, because the resonant process is a complete spin flip which propagates ballistically, the magnetization at intermediate times ($2\sim 6tg$) generally reaches lower values for small $g/J$, and the square seems to start melting more quickly. However, while the small-$g$ dynamics leads to multiple oscillations of magnetization, the large $g/J$ curves monotonously decay to their thermal value. Longer time simulations are required to accurately determine the nature of the thermalized state.
@@
Figure 6: Time evolution of magnetization for the lower right quarter of the square (blue frames) and two rows of surrounding spins (red frames). The y-range of the blue panels is from $-1$ to $1$, while for red panels it is $-1$ to $-0.5$.
@@
Figure 7: Correlations $C_{ij}$ with the $i$-th spin at the corner of the square, at coordinates $(12,12)$ (counted from the upper left corner as $(1,1)$).

### v3

@@ Acknowledgements.
We acknowledge enlightening discussions with Marcello Dalmonte, Roberto Verdel, Simone Notarnicola, Jernej Rudi Finžgar, and Sourav Nandy. The authors acknowledge funding from CN00000013 - Italian Research Center on HPC, Big Data and Quantum Computing (ICSC), the H2020 projects EuRyQa, QuantERA2021 project T-NiSQ, the Quantum Technology Flagship project PASQuanS2, the Italian Departments of Excellence grant 2023-2027 Quantum Frontiers, the German Federal Ministry of Education and Research (BMBF) the funding program quantum technologies, project QRydDemo, and the World Class Research Infrastructure - Quantum Computing and Simulation Center (QCSC) of Padova University. The simulations were performed using the Quantum Green Tea software version 0.3.23 and Quantum Tea Leaves version 0.4.46 [48]. The authors acknowledge computational resources of the HPC center Vega at the Institute of Information Science (IZUM) in Maribor, Slovenia and HPC Leonardo by Cineca, Italy. End matter on perturbative calculations — In the main text (Eq.(2)) we present perturbative expressions for the energies of the first three excitations above the completely polarized initial state. Here, we comment on certain aspects of these result. We Firstly, we find that the predicted excitation energies match the calculated spectral excitations up to a suprisingly large $g/J\sim 1$. This is an indicator that interpreting the dynamics in the confined regime as isolated excitations above the polarized state is accurate. Furthermore, the evolution of the spectrum at larger $g/J$ hints at the underlying reasons for the observed quick thermalization beyond $g/J\sim 2$. $\Delta_{01}$ is the cost of creating a single spin-flip excitation, while $\Delta_{12}$ is the cost of flipping a spin next to an already existing spin flip – growing a spin cluster by one. Quenching to the point where these two process are energetically equally expensive populates not only the first few but also higher excited states, resulting in the formation of larger clusters. At this point, the density of excitations will be large enough for them to overlap and interact, leading to thermalization on a timescale of $1/J$ [35]. Note that aside from the extensive correction common to all states, we do not find a second-order contribution to $E_{2}$. An intuitive interpretation is that in second order of $g$ the corrections arising from coupling $\psi_{2}$ to $\psi_{1}$, and the ones coming from $\psi_{3}$ (superposition of connected three-spin clusters, total domain length of 8), originate from the same physical process. This is the addition or removal of a spin from an existing cluster. Both add two domain walls, and being symmetric in this sense, these contributions cancel out. The processes that couple $\psi_{2}$ to $\psi_{0}$ only appear at fourth order, and we expect that accounting for them would reproduce the curvature of the right and left peaks in Fig. 2(b). Finally, we comment on the differences between the perturbative approach and a similar method of interpreting the spectra – the effective models. Because the dynamics of confined systems is dominated by low energy excitations, it can be accurately described by a simple effective model [16, 14, 24]. This is typically done by projecting the Hamiltonian to a subspace of low-energy excitations (equivalents of our $\psi_{0}$, $\psi_{1}$ and $\psi_{2}$). However, this approach leads to certain inconsistencies. Namely, the matrix element between the ground and first excited state is extensive (indeed, in our case we find $\left\langle\psi_{0}\right|H\left|\psi_{1}\right\rangle=gN$), while others are independent of $N$. Therefore, an effective model which includes $\psi_{0}$ predicts a spectral gap (equivalent of our $\Delta_{01})$ that diverges with system size. To accurately reproduce the excitation energies, the initial vacuum state (equivalent of $\psi_{0}$) should be excluded from the low energy model, and its energy set exactly to zero. Perturbative calculations imply that the origin of the problem is in the arbitrary truncation of the subspace of the effective model. This can indeed be read out from intermediate steps of the calculation, see SM. We find that the extensive corrections come from a state coupling to an extensive number of excitations with one more and one fewer spin flip – it is always possible to flip a decoupled spin in $\mathcal{O}(N^{2})$ sites for an energy penalty of $8J$. The contributions coming from the state with one more and one less spin flip cancel out in the perturbative calculations, and we only find an overall $g^{2}N^{2}/8J$ correction to all states, so that energy differences remain independent of system size. This is because the calculation by design accounts for all existing corrections at a given order of $g$. However, if the Hilbert space is first truncated into an effective model, the coupling between the last retained and the first truncated state is not taken into account. In this case the extensive contributions do not cancel out, and the gaps grow with system size. While the two approaches produce a similar energy diagram, an advantage of the perturbative approach is that it also provides the correction to the ground state energy. It also allows for direct calculations of the eigenstates.
Supplementary Material for: Constrained dynamics and confinement in the two-dimensional quantum Ising model
@@
The numerical calculations presented in this work are performed with tree tensor networks (TTNs) [28, 29], implemented in the QTEA library [48]. In this appendix, we provide qualitative reasoning for the utility of TTNs compared to other tensor networks and technical aspects of the algorithm for time evolution that we apply, and present plots of magnetization and entanglement obtained for a set of maximal bond dimensions.
@@
In TTNs, the state is represented by a tensor network in the shape of a symmetric binary tree, where the bottom tensors (’leaves’ of the tree) carry the physical indices. In comparison to the standard matrix product state (MPS) approach, the tree tensor networks are particularly successful at describing states where distant sites are entangled, such as in chains with long-range interactions. The underlying reason can be understood through a qualitative picture of the way quantum states are encoded in tensor networks. Qualitatively, entanglement between two correlated physical states is transferred through bonds of the network that connect them. The more entanglement there is, the bigger the bond dimension required to accurately describe the state. Generally, this also means entangled states are better represented by tensor networks with a bigger number of links. An extreme example is the projected entangled-pair state (PEPS) ansatz [PEPS], where the tensor network mimics the structure of the 2D lattice and each tensor is directly connected to its nearest neighbours. However, this introduces loops in the network, which leads to technical difficulties. Namely, it is not possible to define the isometry center of a network with loops. Therefore, the algorithms for contracting such networks cannot rely on the underlying symmetries. This leads to poor scaling of calculations of local observables with bond dimension, and limits modern implementations of PEPS to bond dimensions on the order of $\chi\sim 10$.
The TTN ansatz is in between the MPS and PEPS in the sense that it is the best connected tensor network which by construction has no loops. For example, two physical sites separated by $m$ lattice sites in real space are only separated by $\sim\ln m$ links in the TTN and $m$ links in MPS. The entanglement between the sites is thus captured more efficiently than in MPS.
Mapping
An important aspect of TTN simulations in higher dimensions is the mapping of the lattice into a 1D chain, which is then encoded as a TTN. Throughout this work, we use the Hilbert curve as a mapping function. Ref. [30] shows that this is the optimal choice, in the sense that it maximally preserves locality (the sites which are close in the lattice tend to stay close in the mapped 1D chain). We tested various mapping functions, and found that generally the observables converge at smaller bond dimension when using the Hilbert curve.
@@
The most common way to evolve tensor network states in time is with the time-dependent variatonal principle (TDVP) [31, 32]. The most popular is the two-tensor version of the algorithm (TDVP2), where at each step two neighbouring tensors are contracted, evolved in time, and then split and truncated via SVD. This algorithm can grow the bond dimensions between tensors at each time step and is thus a natural choice for simulations starting from weakly entangled states, such as the product states we use. A much faster version of the same concept is the single-tensor TDVP (TDVP1), where each tensor is optimized without contractions with the neighbours. This approach is much cheaper computationally, but cannot grow the bond dimension of the tensors. This presents a problem if the initial state is a product state with $\chi=1$.
However, we find that the calculations are much faster if one pads the initial product state with numerical noise on the order of $10^{-16}$ up to bond dimension $\chi$, and performs the evolution with TDVP1, without truncation. When comparing TDVP2 with initial bond dimension of 1 and maximal of $\chi=100$ to TDVP1 with constant bond dimension of 100, we observe a speedup in computational time of about 30×30$\times$30 × for a $8\times 8$ system. We also tried hybrid approaches (like performing TDVP2 every $n-$th time step and TDVP1 otherwise, with $n$ around 10). The advantage of the hybrid approach is that the initial bond dimension is not $\chi$ but rather gradually increases every $n$-th step. This speeds up the initial time steps, but the advantage is quickly lost once the bond dimension reaches a sizeable fraction of the one used for TDVP1. Overall, we still find the TDVP1 to perform faster.
An advantage might be expected for the easiest cases (small $g/J$ limit) where the bond dimension never reaches the maximal allowed value. However, these cases can also be accurately simulated by a TDVP1 simulation with smaller $\chi$, which is anyway much faster.
@@
This is an advantage on a technical level as well. Having a single convergence parameter ($\chi$) is more desirable compared to having two ($\chi$ and the minimal truncation $\epsilon$) with non-trivial interplay.
@@
In Fig. 1, we show the convergence of (a) magnetization and (c) entanglement with increasing maximal bond dimension $\chi$. The deviation of the magnetization at given $\chi$ with respect to the results obtained with $\chi=160$ are shown in Fig. 1(b).
In calculations with tensor networks it is typically observed that local observables converge first, while global quantities require a much more stringent criteria. Out of the commonly tested quantities, entanglement is typically the last to converge. Convergence of magnetization and of entanglement thus present the easiest and the hardest convergence test, respectively.
Unsuprisingly, we find that the convergence is strongly dependent on the underlying physics. Deep in the confined phase at $g\ll J$, results are well converged for $\chi=60$ or less. At larger $g$, larger $\chi$ is required to obtain well-converged results and the simulations become unreliable much sooner in simulated time.
@@
Supplementary Figure 1: Convergence with bond dimension $\chi$. (a) Magnetization $\sigma^{z}$ of a single spin calculated with increasing bond dimension $\chi$ for a set of $g/J$. (b) Log scale plots of the deviation of magnetization from the results obtained with $\chi=160$, $\delta\left\langle\sigma^{z}\right\rangle=\left|\left\langle\sigma^{z}\right\rangle(\chi)-\left\langle\sigma^{z}\right\rangle(\chi=160)\right|$. (c) Entanglement entropy calculated with increasing bond dimension $\chi$ for a set of small $g/J$.
@@
$H_{0}=-J\sum_{\langle i,j\rangle}\sigma^{z}_{i}\sigma^{z}_{j},$ (S1)
@@
$\delta E^{(1)}=-\left\langle\psi\right|V\left|\psi\right\rangle.$ (S3)
@@
$\begin{split}\delta E^{(2)}&=-\sum_{k\neq\psi}\left\langle\psi\right|V\left|k\right\rangle\frac{1}{E_{k}-E_{0}}\left\langle k\right|V\left|\psi\right\rangle\\ &=-\sum_{k\neq\psi}\frac{1}{E_{k}-E_{0}}\left|\left\langle\psi\right|V\left|k\right\rangle\right|^{2},\end{split}$ (S4)
@@
$\begin{split}\delta E_{0}^{(2)}&=-\sum_{k}\left\langle\psi_{0}\right|V\left|k\right\rangle\frac{1}{E_{k}-E_{0}}\left\langle k\right|V\left|\psi_{0}\right\rangle\\ &=-\sum_{k}\frac{1}{E_{k}-E_{0}}\left|\left\langle\psi_{0}\right|V\left|k\right\rangle\right|^{2}.\end{split}$ (S5)
@@
The sum in Eq. (S5) thus gives
$\delta E_{0}^{(2)}=-\sum_{i}^{N}\frac{g^{2}}{8J}=-\frac{g^{2}}{8J}N^{2}.$ (S6)
@@
$\left|\psi_{1}\right\rangle=\frac{1}{\sqrt{N^{2}}}\sum_{i}\sigma^{x}_{i}\left|\psi_{0}\right\rangle.$ (S7)
@@
$\delta E_{1}^{(2)}=\sum_{k}\frac{1}{E_{k}-E_{1}}\left|\left\langle\psi_{1}\right|V\left|k\right\rangle\right|^{2},$ (S8)
@@
$\left|k\right\rangle=\sigma^{x}_{i}\sigma^{x}_{j}\left|\psi_{0}\right\rangle.$ (S9)
@@
$\begin{split}\left\langle\psi_{0}\right|V\left|\psi_{1}\right\rangle&=\left\langle\psi_{0}\right|-g\sum_{i}\sigma^{x}_{i}\frac{1}{\sqrt{N^{2}}}\sum_{j}\sigma^{x}_{j}\left|\psi_{0}\right\rangle\\ &=-\frac{g}{\sqrt{N^{2}}}\sum_{ij}\left\langle\psi_{0}\right|\sigma^{x}_{i}\sigma^{x}_{j}\left|\psi_{0}\right\rangle\\ &=-\frac{g}{\sqrt{N^{2}}}\sum_{ij}\delta_{ij}\\ &=-\frac{g}{\sqrt{N^{2}}}N^{2},\end{split}$ (S10)
@@
$\frac{1}{E_{0}-E_{1}}\left|\left\langle\psi_{0}\right|V\left|\psi_{1}\right\rangle\right|^{2}=-\frac{1}{8J}\times g^{2}N^{2}.$ (S11)
@@
$\begin{split}\left\langle\psi_{2}^{\prime}\right|_{kl}V\left|\psi_{1}\right\rangle&=-\frac{g}{\sqrt{N^{2}}}\left\langle\psi_{2}^{\prime}\right|_{kl}\sum_{ij}\sigma^{x}_{i}\sigma^{x}_{j}\left|\psi_{0}\right\rangle\\ &=-\frac{g}{\sqrt{N^{2}}}\times 2,\end{split}$ (S12)
@@
$2N^{2}\frac{1}{E_{2}^{\prime}-E_{1}}\left|\left\langle\psi_{2}^{\prime}\right|_{kl}V\left|\psi_{1}\right\rangle\right|^{2}=2N^{2}\times\frac{4g^{2}}{N^{2}}\times\frac{1}{4J}=\frac{2g^{2}}{J}.$ (S13)
@@
$\left\langle\psi_{2}^{\prime\prime}\right|_{kl}V\left|\psi_{1}\right\rangle=-\frac{g}{\sqrt{N^{2}}}\times 2,$ (S14)
@@
$\frac{1}{2}N^{2}(N^{2}-5)\times\frac{4g^{2}}{N^{2}}\times\frac{1}{8J}=2\frac{g^{2}}{8J}\left(N^{2}-5\right).$ (S15)
@@
$\begin{split}\delta E_{1}^{(2)}&=-\left(-\frac{g^{2}}{8J}N^{2}+\frac{2g^{2}}{J}+\frac{g^{2}}{8J}2\left(N^{2}-5\right)\right)\\ &=-\frac{g^{2}}{8J}\left(N^{2}+6\right).\end{split}$ (S16)
Importantly, the extensive ($\propto N^{2}$) contribution is the same as for the ground state, see Eq. (S6), resulting in a $N$-independent gap.
@@
$\left|\psi_{2}^{\prime}\right\rangle=\frac{1}{2N^{2}}\sum_{ij}\sigma^{x}_{i}\sigma^{x}_{j}\left|\psi_{0}\right\rangle,$ (S17)
@@
$\left\langle\psi_{1}\right|V\left|\psi_{2}^{\prime}\right\rangle=\left\langle\psi_{0}\right|\sigma^{x}_{i}\sum_{j}g\sigma^{x}_{j}\frac{1}{\sqrt{2N^{2}}}\sum_{kl}\sigma^{x}_{k}\sigma^{x}_{l}\left|\psi_{0}\right\rangle=g\frac{4}{\sqrt{2N^{2}}}.$ (S18)
@@
$N^{2}\times\frac{-1}{4J}\times\frac{16g^{2}}{2N^{2}}.$ (S19)
@@
$\begin{split}\left\langle\psi_{3}^{\prime}\right|V\left|\psi_{2}^{\prime}\right\rangle=g\frac{2}{\sqrt{2N^{2}}}.\end{split}$ (S20)
@@
$6N^{2}\times\frac{1}{4J}\times\frac{4g^{2}}{2N^{2}}.$ (S21)
@@
$\left\langle\psi_{3}^{\prime\prime}\right|V\left|\psi_{2}^{\prime}\right\rangle=g\frac{1}{2N^{2}},$ (S22)
@@
$N^{2}\left(N^{2}-8\right)\times\frac{1}{4J}\times\frac{g^{2}}{2N^{2}}.$ (S23)
@@
$\delta E_{2}^{(2)}=-g\frac{g^{2}}{8J}N^{2}.$ (S24)
See the main text around main text Fig. 2 for a qualitative interpretation of the obtained result.
Additional results for the interface
In this section we provide additional results of the time evolution of magnetization in the case of an interface. These are obtained in a $16\times 8$ system on a cylinder (periodic boundary conditions along the shorter and open boundary along the longer dimension).
As described in the main text, we find three dynamical regimesl, and smooth transitions between them. For small $g/J$, the system is constrained. The spins at the interface oscillate via the edge mode. At intermediate $g/J\sim 1--1.7$, the dominant process is the interface broadening, while the rest of the system remains static. At large $g/J$ the dynamical constraint does not play a role. All spins oscillate simultaneously with a frequency of approximately $J$, and the system thermalizes quickly.
Supplementary Figure 2: Time evolution of the magnetization $\sigma_{z}$ for a stripe initial state; $16\times 8$ system, initialized with two domains. The interface is between the $7$th and $8$th spin in the longer dimension. (a) Cuts of magnetization for an extended set of $g/J$. (b) Magnetization for five spins, one in the $\uparrow$ polarized domain, and four in the $\downarrow$ polarized domain. The spin coordinates are denoted in the brackets in each panel.
@@
In this section, we provide additional results for the dynamics of the square. Fig. 3 shows a grid of magnetization for the lower right quadrant of the square (the panels with blue frames) and two surrounding rows and columns (red frames). Time is rescaled by $g$, as in main text Fig. 5.
As discussed in the main text, we observe a crossover between two limiting behaviours; at small $g/J$ the dynamics consists of a sequence of resonant spin flips which originate from the corner, while with increasing $g/J$ the spins at the interface melt simultaneously. Spins within the square remain inert until the wave of melting reaches them. Interestingly, because the resonant process is a complete spin flip which propagates ballistically, the magnetization at intermediate times ($\sim 6tg$) generally reaches lower values for small $g/J$, and the square seems to start melting more quickly. However, while the small-$g$ dynamics leads to multiple oscillations of magnetization, the large $g/J$ curves monotonously decay to their thermal value. Longer time simulations are required to accurately determine the nature of the thermalized state.
@@
Supplementary Figure 3: Time evolution of magnetization for the lower right quarter of the square (blue frames) and two rows of surrounding spins (red frames). The y-range of the blue panels is from $-1$ to $1$, while for red panels it is $-1$ to $-0.5$.
@@
Supplementary Figure 4: Correlations $C_{ij}$ with the $i$-th spin at the corner of the square, at coordinates $(12,12)$ (counted from the upper left corner as $(1,1)$).